# Fractal counterparts of density combinatorics theorems in continued fractions

### v1

@@ 3.2. Adding digits into seed sets
Let $M_{0}=0$, and let $(M_{k})_{k=1}^{\infty}$ be a strictly increasing sequence of positive integers. Let $(W_{k})_{k=1}^{\infty}$ be a sequence of finite subsets of $A$. Let $y\in R_{t,L}(B)$ and put $a_{0}(y)=0$ for convenience. For each $k\geq 1$ with $W_{k}\neq\emptyset$, we add the elements of $W_{k}$ into the digit sequence $(a_{n}(y))_{n=1}^{\infty}$ in the increasing order [^3] This ordering in $W_{k}$ is not essential. For a related comment, see the footnote in Corollary 1.3. to define a new sequence
@@
where $W_{k}=\{w^{(k)}_{i}\colon i=1,\ldots,i(k)\}$, $w^{(k)}_{1}<\cdots<w^{(k)}_{i(k)}$. Let $x(y)$ denote the point in $(0,1)\setminus\mathbb{Q}$ whose regular continued fraction expansion (1.1) is given by this new sequence. Let $R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$ denote the collection of these points:
@@
denote the inverse of this map. We call $f_{t,L,B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty}}$ an elimination map. It eliminates all the digits in $\bigcup_{k=1}^{\infty}W_{k}$. Clearly, if $A\cap B=\emptyset$ and the elements of $(W_{k})_{k=1}^{\infty}$ are pairwise disjoint then $R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$ is contained in $E$.
3.3. Almost Lipschitzness of the elimination map
We need to deduce a lower bound of the Hausdorff dimension of the set $R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$ constructed in §$\S$3.2. This can be done by evaluating the Hölder exponent of the elimination map. This map is merely Hölder in general and as a result the Hausdorff dimension of $R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$ may drop from that of $R_{t,L}(B)$. However, under appropriate choices of $(M_{k})_{k=1}^{\infty}$ and $(W_{k})_{k=1}^{\infty}$, the Hölder exponent can be made closer and closer to $1$ on deeper levels of $R_{t,L}(B)$, so that the elimination map becomes almost Lipschitz, as stated in the proposition below.
@@ 3.3. Almost Lipschitzness of the elimination map
Let $(\varepsilon_{k})_{k=1}^{\infty}$ be a monotone decreasing sequence of positive reals converging to 00. Let $A$, $B$ be infinite subsets of $\mathbb{N}$ and suppose $B$ does not contain consecutive integers. Let $t>L>1$, $t>2$ and suppose $R_{t,L}(B)\neq\emptyset$. Let $M_{0}=0$ and let $(M_{k})_{k=1}^{\infty}$ be a strictly increasing sequence of positive integers, and let $(W_{k})_{k=1}^{\infty}$ be a sequence of finite subsets of $A$ such that for every $k\geq 1$,
@@
Then $f=f_{t,L,B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty}}$ is almost Lipschitz.
Proof.
A key observation motivating the proof of this proposition is that the influence of adding the elements of $W_{k}$ is absorbed into the contribution from the digits $a_{M_{k-1}+1}(y),\ldots,a_{M_{k}}(y)$, $y\in R_{t,L}(B)$ by virtue of (3.10). The role of (3.11) is auxiliary, to be used after (3.10).
@@
We estimate $|x_{1}-x_{2}|$ from below and $|y_{1}-y_{2}|$ from above. Using the first inequality in Lemma 2.1 and $a_{i}(y_{1})\leq Lt^{i}$ for $11\leq i\leq n+$ 1 we have
@@
Setting $C_{k}=D_{k}^{\frac{1}{1+4\varepsilon_{k}}}$ yields $|f(x_{1})-f(x_{2})|\leq C_{k}|x_{1}-x_{2}|^{\frac{1}{1+4\varepsilon_{k}}}$. Since $\varepsilon_{k}\to 0$ as $k\to\infty$, $f$ is almost Lipschitz. ∎
@@ 4. Proofs of the main results
In this last section we complete the proofs of the main results. Since Theorem 1.1(a) follows from Theorem 1.4, we prove Theorem 1.4 in §$\S$4.1, and then prove Theorem 1.1(b) in §$\S$4.2. Finally we prove Theorem 1.13 in §$\S$4.3.
@@
Let $S\subset\mathbb{N}$ have polynomial density with exponent $\alpha$. By Proposition 3.1 there exist $S_{*}\subset S$ not containing consecutive integers and positive reals $t$, $L$ with $t>L>1$ and $t>2$ such that $R_{t,L}(S_{*})$ is an extreme seed set of Hausdorff dimension $1/(2\alpha)$. Let $A\subset S$ satisfy $\overline{d}(A|S)>0$. We have
@@
Let $(\varepsilon_{k})_{k=1}^{\infty}$ be a monotone decreasing sequence of positive reals converging to 00. We set $M_{0}=0$ for convenience, and choose a strictly increasing sequence $(M_{k})_{k=1}^{\infty}$ of positive integers greater than one such that all the following conditions hold:
@@
$\frac{1}{(\sqrt{M_{k}}+Lt^{M_{k-1}}+1)^{2\sqrt{M_{k}}}}\geq\frac{1}{t^{2\varepsilon_{k}M_{k}^{2}}}\geq\left(\prod_{i=M_{k-1}+1}^{M_{k}}\frac{1}{t^{2i}}\right)^{3\varepsilon_{k}}\ \text{ for every }k\geq 1;$ (4.5)
@@
Since $S$ has polynomial density, it intersects both $I_{k}$ and $[1,\sqrt{M_{k}}]$ for all sufficiently large $k$. In the next two paragraphs we show that the two differences
#$\frac{\#(A\cap I_{k})}{\#(S\cap I_{k})}-\frac{\#W_{k}}{\#(S\cap I_{k})}=\frac{\#(S_{*}\cap I_{k})}{\#(S\cap I_{k})}$
@@
|#$\begin{split}\left|\frac{\#W_{k}}{\#(S\cap I_{k})}-\frac{\#((A\setminus S_{*})\cap[1,\sqrt{M_{k}}])}{\#(S\cap[1,\sqrt{M_{k}}])}\right|\leq&\left|\frac{\#W_{k}-\#((A\setminus S_{*})\cap[1,\sqrt{M_{k}}])}{\#(S\cap I_{k})}\right|\\ &+\left|\frac{\#(S\cap[1,\sqrt{M_{k}}])-\#(S\cap I_{k})}{\#(S\cap I_{k})}\right|\end{split}$
converge to 00 as $k\to\infty$.
Since $S$ has polynomial decay with exponent $\alpha\geq 1$, there exist constants $C_{1}>0$ and $\beta\geq 0$ such that
@@
Since $S$ has polynomial decay with exponent $\alpha\geq 1$ and $\overline{d}(S_{*}|S)=0$, there exist $C_{2}>0$ and for any $\varepsilon>0$ there exists $k_{0}\geq 1$ such that for every $k\geq k_{0}$,
#$\#(S_{*}\cap I_{k})\leq\#(S_{*}\cap[1,Lt^{M_{k-1}}+\sqrt{M_{k}}]\leq\varepsilon\cdot\frac{C_{2}(Lt^{M_{k-1}}+\sqrt{M_{k}})^{1/\alpha}}{(\log(Lt^{M_{k-1}}+\sqrt{M_{k}}))^{\beta}}.$
@@
|#$\begin{split}&|\#W_{k}-\#((A\setminus S_{*})\cap[1,\sqrt{M_{k}}])|\\ =&|\#((A\setminus S_{*})\cap[Lt^{M_{k-1}}+1,\sqrt{M_{k}}])+\#((A\setminus S_{*})\cap[\sqrt{M_{k}}+1,\sqrt{M_{k}}+Lt^{M_{k-1}}])\\ &-\#((A\setminus S_{*})\cap[1,Lt^{M_{k-1}}])-\#((A\setminus S_{*})\cap[Lt^{M_{k-1}}+1,\sqrt{M_{k}}])|\leq Lt^{M_{k-1}}.\end{split}$
@@
$E_{S,A}=R_{t,L}(S_{*},A\setminus S_{*},(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty}).$
The construction of this set was described in §$\S$3.2. For each $y\in R_{t,L}(S_{*})$ we add the elements of $W_{k}$ into the digit sequence $(a_{n}(y))_{n=1}^{\infty}$ at position $M_{k}$ to define a new sequence. Then $E_{S,A}$ is the set of points whose regular continued fraction expansions are given by these new sequences.
We claim that $E_{S,A}\subset E$. Indeed, let $x\in E_{S,A}$ and let $y$ be the corresponding point in $R_{t,L}(S_{*})$, namely $y=f(x)$. Since $(a_{n}(y))_{n=1}^{\infty}$ is a strictly increasing sequence in $S_{*}$ and $\bigsqcup_{k=1}^{\infty}W_{k}\subset A\setminus S_{*}$, for all $m$, $n\in\mathbb{N}$ distinct we have $a_{m}(x)\neq a_{n}(x)$, namely $x\in E$. By the construction of $E_{S,A}$, we have
#$\#W_{k}\leq\#\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\cap I_{k}\leq\#(A\cap I_{k}),$
@@
$\lim_{k\to\infty}\frac{\#\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\cap I_{k}}{\#(S\cap I_{k})}=\overline{d}(A|S).$ (4.11)
@@
It is clear from the construction that $E_{S,A}\subset\{x\in E\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}$. Combining (4.12) and Corollary 3.4 we obtain (1.6).
Since $S$ has polynomial density with exponent $\alpha$, there exist $C_{1},C_{2}>0$ and $\beta\geq 0$ such that for all sufficiently large $k\geq 1$ we have
#$\frac{\#(S\cap[1,\min I_{k}))}{\#(S\cap[1,\max I_{k}])}\leq C_{1}C_{2}^{-1}\frac{(\min I_{k})^{1/\alpha}}{(\max I_{k})^{1/\alpha}}\frac{(\log(\max I_{k}))^{\beta}}{(\log(\min I_{k}))^{\beta}}.$
@@
$\lim_{k\to\infty}\frac{\#(S\cap I_{k})}{\#(S\cap[1,\max I_{k}])}=1-\lim_{k\to\infty}\frac{\#(S\cap[1,\min I_{k}])}{\#(S\cap[1,\max I_{k}])}=1.$
@@
$\begin{split}\overline{d}\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A|S\right)\geq&\limsup_{k\to\infty}\frac{\#\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\cap[1,\max I_{k}]}{\#(S\cap[1,\max I_{k}])}\\ \geq&\limsup_{k\to\infty}\frac{\#\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\cap I_{k}}{\#(S\cap[1,\max I_{k}])}\\ =&\lim_{k\to\infty}\frac{\#(S\cap I_{k})}{\#(S\cap[1,\max I_{k}])}\\ &\times\lim_{k\to\infty}\frac{\#\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\cap I_{k}}{\#(S\cap I_{k})}\\ =&\overline{d}(A|S).\end{split}$
@@
In principle we proceed along the line of §$\S$4.1, but there are a number of technical differences. Let $S\subset\mathbb{N}$ satisfy $\overline{d}(S)=0$ and $d_{\rm B}(S)>0$. We set $M_{0}=N_{0}=0$ for convenience, and take two sequences $(M_{k})_{k=1}^{\infty}$, $(N_{k})_{k=1}^{\infty}$ of positive integers such that $(N_{k})_{k=1}^{\infty}$ is strictly increasing and
$\lim_{k\to\infty}\frac{1}{N_{k}}\#({S\cap[M_{k},M_{k}+N_{k}]})=d_{\rm B}(S).$ (4.13)
If $\sup_{k\geq 1}M_{k}<\infty$ then we have $d_{\rm B}(S)=\overline{d}(S)$. This case is covered by Theorem 1.1(a). In what follows we assume $\sup_{k\geq 1}M_{k}=\infty$. Taking subsequences of $(M_{k})_{k=1}^{\infty}$ and $(N_{k})_{k=1}^{\infty}$ if necessary, we may assume $(M_{k})_{k=1}^{\infty}$ is strictly increasing and $M_{k}+N_{k}<M_{k+1}$ for every $k\geq 1$. For each $k\geq 1$, set
$I_{k}=[M_{k-1},M_{k-1}+N_{k-1}].$
@@
$\lim_{k\to\infty}\frac{\#(S\cap I_{k})}{\#(\mathbb{N}\cap I_{k})}=d_{\rm B}(S).$ (4.14)
Let $(\varepsilon_{k})_{k=1}^{\infty}$ be a monotone decreasing sequence of positive reals converging to 00. Taking subsequences of $(M_{k})_{k=1}^{\infty}$ and $(N_{k})_{k=1}^{\infty}$ if necessary, we may assume
@@
$\prod_{i\in S\cap I_{k}}\frac{1}{(i+1)^{2}}\geq\left(\prod_{i=M_{k-1}+1}^{M_{k}}\frac{1}{t^{2i}}\right)^{3\varepsilon_{k}}\ \text{ for every }k\geq 1\text{ with }S\cap I_{k}\neq\emptyset.$ (4.17)
Since $\overline{d}(S)=0$, the set $\mathbb{N}\setminus S$ has polynomial density with $\alpha=1$, $\beta=0$. By Proposition 3.1, there exist a subset $(\mathbb{N}\setminus S)_{*}$ of $\mathbb{N}\setminus S$ not containing consecutive integers and positive reals $t,L$ with $t>L>1$ and $t>2$ such that $R_{t,L}((\mathbb{N}\setminus S)_{*})$ is an extreme seed set of Hausdorff dimension $1/2$. We set
$F_{S}=R_{t,L}((\mathbb{N}\setminus S)_{*},S,(M_{k})_{k=1}^{\infty},(S\cap I_{k})_{k=1}^{\infty}).$
The construction of this set was described in §$\S$3.2. For each $y\in R_{t,L}((\mathbb{N}\setminus S)_{*})$ we add the elements of $S\cap I_{k}$ into the digit sequence $(a_{n}(y))_{n=1}^{\infty}$ at position $M_{k}$ to define a new sequence. Then $F_{S}$ is the set of points whose regular continued fraction expansions are given by these new sequences.
@@
$\dim_{\rm H}F_{S}\geq\dim_{\rm H}R_{t,L}((\mathbb{N}\setminus S)_{*})=\frac{1}{2}.$ (4.18)
@@
$\bigcup_{n\in\mathbb{N}}\bigcap_{x\in F_{S}}\{a_{n}(x)\}\cap S\cap I_{k}=S\cap I_{k}.$
@@
$\begin{split}d_{\rm B}\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in F_{S}}\{a_{n}(x)\}\cap S\right)&\geq\limsup_{k\to\infty}\frac{\#\bigcup_{n\in\mathbb{N}}\bigcap_{x\in F_{S}}\{a_{n}(x)\}\cap S\cap I_{k}}{\#(\mathbb{N}\cap I_{k})}\\ &=\lim_{k\to\infty}\frac{\#(S\cap I_{k})}{\#(\mathbb{N}\cap I_{k})}=d_{\rm B}(S).\end{split}$

### v2

@@ 3.4. Adding digits into seed sets (Step 3)
Let $M_{0}=0$, and let $(M_{k})_{k=1}^{\infty}$ be a strictly increasing sequence of positive integers. Let $(W_{k})_{k=1}^{\infty}$ be a sequence of finite subsets of $A$. Let $y\in R_{t,L}(B)$ and put $a_{0}(y)=0$ for convenience. For each $k\geq 1$ with $W_{k}\neq\emptyset$, we add the elements of $W_{k}$ into the digit sequence $(a_{n}(y))_{n=1}^{\infty}$ in the increasing order to define a new sequence
@@
where
$W_{k}=\{w^{(k)}_{i}\colon i=1,\ldots,i(k)\},\ w^{(k)}_{1}<\cdots<w^{(k)}_{i(k)}.$
Let $x(y)$ denote the point in $(0,1)\setminus\mathbb{Q}$ whose regular continued fraction expansion (1.1) is given by this new sequence. Let $R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$ denote the collection of these points:
@@
denote the inverse of this map. We call $f_{t,L,B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty}}$ an elimination map. Clearly, if $A\cap B=\emptyset$ and the elements of $(W_{k})_{k=1}^{\infty}$ are pairwise disjoint then $R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$ is contained in $E$. Moreover, the elimination map $f$ sends $x\in R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$ to $f(x)\in R_{t,L}(B)$ whose digit sequence is given by eliminating elements of $\bigcup_{k=1}^{\infty}W_{k}$ from the digit sequence of $x$.
3.5. Almost Lipschitzness of the elimination map (Step 3)
We need to deduce a lower bound of the Hausdorff dimension of the set $R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$ constructed in Section 3.4. This can be done by evaluating the Hölder exponent of the elimination map. This map is merely Hölder in general and as a result the Hausdorff dimension of $R_{t,L}(B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$ may drop from that of $R_{t,L}(B)$. However, under appropriate choices of $(M_{k})_{k=1}^{\infty}$ and $(W_{k})_{k=1}^{\infty}$, the Hölder exponent can be made closer and closer to $1$ on deeper levels of $R_{t,L}(B)$, so that the elimination map becomes almost Lipschitz, as stated in the proposition below.
@@ 3.5. Almost Lipschitzness of the elimination map (Step 3)
Let $(\varepsilon_{k})_{k=1}^{\infty}$ be a decreasing sequence of positive reals converging to 0. Let $A$, $B$ be infinite subsets of $\mathbb{N}$ and suppose $B$ does not contain consecutive integers. Let $t>L>1$, $t>2$ and suppose $R_{t,L}(B)\neq\emptyset$. Let $M_{0}=0$ and let $(M_{k})_{k=1}^{\infty}$ be a strictly increasing sequence of positive integers, and let $(W_{k})_{k=1}^{\infty}$ be a sequence of finite subsets of $A$ such that for every $k\geq 1$,
@@
Then the elimination map $f=f_{t,L,B,A,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty}}$ is almost Lipschitz.
Regarding a proof of this proposition, a key observation is that the influence of adding the elements of $W_{k}$ is absorbed into the contribution from the digits $a_{M_{k-1}+1}(y),\ldots,a_{M_{k}}(y)$, $y\in R_{t,L}(B)$ by virtue of (3.11). The role of (3.12) is auxiliary, to be used after (3.11) is.
Proof of Proposition 3.5.
@@
We estimate $|x_{1}-x_{2}|$ from below and $|y_{1}-y_{2}|$ from above. Using the first inequality in Lemma 2.1 and $a_{i}(y_{1})\leq Lt^{i}$ for $1\leq i\leq n+1$, we have
@@
Setting $C_{k}=D_{k}^{\frac{1}{1+4\varepsilon_{k}}}$ yields $|f(x_{1})-f(x_{2})|\leq C_{k}|x_{1}-x_{2}|^{\frac{1}{1+4\varepsilon_{k}}}$. Since $\varepsilon_{k}\to 0$ as $k\to\infty$, $f$ is almost Lipschitz. The proof of Proposition 3.5 is complete.∎
@@ 4. Proofs of the main results
In this section we complete the proofs of our main results, following the three steps developed in Section 3. In Section 4.1 we prove Theorem 1.4, which implies Theorem 1.1(a). We prove Theorem 1.1(b) in Section 4.2. Central to these proofs is to verify the key condition (3.11). To this end, we carefully choose positions of digits to be added in Step 3, taking advantage of the non-monotonicity of digits for points in $E\setminus J$. Finally in Section 4.3 we prove Theorem 1.13.
@@
Let $S\subset\mathbb{N}$ have polynomial density with exponent $\alpha\geq 1$. By Proposition 3.1, there exist $S_{*}\subset S$ not containing consecutive integers and positive reals $t$, $L$ with $t>L>1$ and $t>2$ such that $R_{t,L}(S_{*})$ is an extreme seed set of Hausdorff dimension $1/(2\alpha)$. Let $A\subset S$ satisfy $\overline{d}(A|S)>0$. We have
@@
Let $(\varepsilon_{k})_{k=1}^{\infty}$ be a decreasing sequence of positive reals converging to 0. We set $M_{0}=0$ for convenience, and choose a strictly increasing sequence $(M_{k})_{k=1}^{\infty}$ of positive integers greater than one such that all the following conditions hold:
@@
$\frac{1}{(Lt^{M_{k-1}}+\sqrt{M_{k}}+1)^{2\sqrt{M_{k}}}}\geq\frac{1}{t^{2\varepsilon_{k}M_{k}^{2}}}\geq\left(\prod_{i=M_{k-1}+1}^{M_{k}}\frac{1}{t^{2i}}\right)^{3\varepsilon_{k}}\ \text{ for every }k\geq 1;$ (4.5)
@@
Since $S$ has polynomial density, it intersects both $I_{k}$ and $1,\sqrt{M_{k}}]$ for all sufficiently large $k$. In the next two paragraphs we show that the two differences
$\frac{\#(A\cap I_{k})}{\#(S\cap I_{k})}-\frac{\#W_{k}}{\#(S\cap I_{k})}=\frac{\#(S_{*}\cap I_{k})}{\#(S\cap I_{k})}$
@@
$\begin{split}\left|\frac{\#W_{k}}{\#(S\cap I_{k})}-\frac{\#((A\setminus S_{*})\cap[1,\sqrt{M_{k}}])}{\#(S\cap[1,\sqrt{M_{k}}])}\right|\leq&\left|\frac{\#W_{k}-\#((A\setminus S_{*})\cap[1,\sqrt{M_{k}}])}{\#(S\cap I_{k})}\right|\\ &+\left|\frac{\#(S\cap[1,\sqrt{M_{k}}])-\#(S\cap I_{k})}{\#(S\cap I_{k})}\right|\end{split}$
converge to 0 as $k\to\infty$.
Since $S$ has polynomial density with exponent $\alpha\geq 1$, there exist constants $C_{1}>0$ and $\beta\geq 0$ such that
@@
Since $S$ has polynomial density with exponent $\alpha\geq 1$ and $\overline{d}(S_{*}|S)=0$, there exist $C_{2}>0$ and for any $\varepsilon>0$ there exists $k_{0}\geq 1$ such that for every $k\geq k_{0}$,
$(S_{*}\cap I_{k})\leq\#(S_{*}\cap[1,Lt^{M_{k-1}}+\sqrt{M_{k}}])\leq\varepsilon\cdot\frac{C_{2}(Lt^{M_{k-1}}+\sqrt{M_{k}})^{1/\alpha}}{(\log(Lt^{M_{k-1}}+\sqrt{M_{k}}))^{\beta}}.$
@@
Figure 2. The white and black dots altogether indicate part of the graph of the map $n\in\mathbb{N}\mapsto a_{n}(x)\in\mathbb{N}$ for each $x\in E_{S,A}=R_{t,L}(S_{*},A\setminus S_{*},(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty})$. The black dots indicate the digits in $W_{k}$ (resp. $W_{k+1}$) inserted in between time $M_{k}$ and time $M_{k}+1$ (resp. between time $M_{k+1}$ and time $M_{k+1}+1$).
@@
$\begin{split}&|\#W_{k}-\#((A\setminus S_{*})\cap[1,\sqrt{M_{k}}])|\\ =&|\#((A\setminus S_{*})\cap[Lt^{M_{k-1}}+1,\sqrt{M_{k}}])+\#((A\setminus S_{*})\cap[\sqrt{M_{k}}+1,\sqrt{M_{k}}+Lt^{M_{k-1}}])\\ &-\#((A\setminus S_{*})\cap[1,Lt^{M_{k-1}}])-\#((A\setminus S_{*})\cap[Lt^{M_{k-1}}+1,\sqrt{M_{k}}])|\leq Lt^{M_{k-1}}.\end{split}$
@@
$E_{S,A}=R_{t,L}(S_{*},A\setminus S_{*},(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty}).$ (4.11)
The construction of this set was described in Section 3.4. For each $y\in R_{t,L}(S_{*})$ we add the elements of $W_{k}$ into the digit sequence $(a_{n}(y))_{n=1}^{\infty}$ at position $M_{k}$ to define a new sequence. Then $E_{S,A}$ is the set of points whose regular continued fraction expansions are given by these new sequences. It is clear from the construction that for any $x\in E_{S,A}$, the digit sequence $\{a_{n}(x)\}_{n\in\mathbb{N}}$ is not monotone increasing, as shown in Figure 2.
We claim that $E_{S,A}\subset E$. Indeed, let $x\in E_{S,A}$ and let $y$ be the corresponding point in $R_{t,L}(S_{*})$, namely $y=f(x)$. Since $(a_{n}(y))_{n=1}^{\infty}$ is a strictly increasing sequence in $S_{*}$ and $\bigsqcup_{k=1}^{\infty}W_{k}\subset A\setminus S_{*}$, for all $m$, $n\in\mathbb{N}$ with $m\neq n$ we have $a_{m}(x)\neq a_{n}(x)$, namely $x\in E$. By the construction of $E_{S,A}$, we have
$W_{k}\leq\#\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\cap I_{k}\right)\leq\#(A\cap I_{k}),$
@@
$\lim_{k\to\infty}\frac{\#\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\cap I_{k}\right)}{\#(S\cap I_{k})}=\overline{d}(A|S).$ (4.12)
@@
It is clear from the construction that $E_{S,A}\subset\{x\in E\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}$. Combining (4.13) and Proposition 3.4 we obtain (1.6).
Since $S$ has polynomial density with exponent $\alpha\geq 1$, there exist $C_{1},C_{2}>0$ and $\beta\geq 0$ such that for all sufficiently large $k\geq 1$ we have
$\frac{\#(S\cap[1,\min I_{k}))}{\#(S\cap[1,\max I_{k}])}\leq C_{1}^{-1}C_{2}\frac{(\min I_{k})^{1/\alpha}}{(\max I_{k})^{1/\alpha}}\frac{(\log(\max I_{k}))^{\beta}}{(\log(\min I_{k}))^{\beta}}.$
@@
$\lim_{k\to\infty}\frac{\#(S\cap I_{k})}{\#(S\cap[1,\max I_{k}])}=1-\lim_{k\to\infty}\frac{\#(S\cap[1,\min I_{k}))}{\#(S\cap[1,\max I_{k}])}=1.$
@@
$\begin{split}\overline{d}\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A|S\right)\geq&\limsup_{k\to\infty}\frac{\#\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\cap[1,\max I_{k}]\right)}{\#(S\cap[1,\max I_{k}])}\\ \geq&\limsup_{k\to\infty}\frac{\#\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\cap I_{k}\right)}{\#(S\cap[1,\max I_{k}])}\\ =&\lim_{k\to\infty}\frac{\#(S\cap I_{k})}{\#(S\cap[1,\max I_{k}])}\\ &\times\lim_{k\to\infty}\frac{\#\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S,A}}\{a_{n}(x)\}\cap A\cap I_{k}\right)}{\#(S\cap I_{k})}\\ =&\overline{d}(A|S).\end{split}$
@@
In principle we proceed along the line of Section 4.1, but there are a number of technical differences. Let $S\subset\mathbb{N}$ satisfy $\overline{d}(S)=0$ and $d_{\rm B}(S)>0$. We set $M_{0}=N_{0}=0$ for convenience, and take two sequences $(M_{k})_{k=1}^{\infty}$, $(N_{k})_{k=1}^{\infty}$ of positive integers such that $(N_{k})_{k=1}^{\infty}$ is strictly increasing and
$\lim_{k\to\infty}\frac{\#({S\cap[M_{k},M_{k}+N_{k}]})}{N_{k}}=d_{\rm B}(S).$
If $\sup_{k\geq 1}M_{k}<\infty$ then we have $d_{\rm B}(S)=\overline{d}(S)$. This case is covered by Theorem 1.1(a). In what follows we assume $\sup_{k\geq 1}M_{k}=\infty$. Taking subsequences of $(M_{k})_{k=1}^{\infty}$ and $(N_{k})_{k=1}^{\infty}$ if necessary, we may assume $(M_{k})_{k=1}^{\infty}$ is strictly increasing and $M_{k}+N_{k}<M_{k+1}$ for every $k\geq 1$. For each $k\in\mathbb{N}$ we put
$I_{k}=[M_{k-1},M_{k-1}+N_{k-1}]$
and
$W_{k}=S\cap I_{k}.$
@@
$\lim_{k\to\infty}\frac{\#W_{k}}{\#(\mathbb{N}\cap I_{k})}=d_{\rm B}(S).$ (4.14)
Let $(\varepsilon_{k})_{k=1}^{\infty}$ be a decreasing sequence of positive reals converging to 0. Taking subsequences of $(M_{k})_{k=1}^{\infty}$ and $(N_{k})_{k=1}^{\infty}$ if necessary, we may assume
@@
$\prod_{i\in W_{k}}\frac{1}{(i+1)^{2}}\geq\left(\prod_{i=M_{k-1}+1}^{M_{k}}\frac{1}{t^{2i}}\right)^{3\varepsilon_{k}}\ \text{ for every }k\geq 1\text{ with }S\cap I_{k}\neq\emptyset.$ (4.17)
Since $\overline{d}(S)=0$, the set $\mathbb{N}\setminus S$ has polynomial density with $\alpha=1$, $\beta=0$. By Proposition 3.1, there exist a subset $(\mathbb{N}\setminus S)_{*}$ of $\mathbb{N}\setminus S$ not containing consecutive integers, and positive reals $t,L$ with $t>L>1$ and $t>2$ such that $R_{t,L}((\mathbb{N}\setminus S)_{*})$ is an extreme seed set of Hausdorff dimension $1/2$. We set
$F_{S}=R_{t,L}((\mathbb{N}\setminus S)_{*},S,(M_{k})_{k=1}^{\infty},(W_{k})_{k=1}^{\infty}).$ (4.18)
The construction of this set was described in Section 3.4. For each $y\in R_{t,L}((\mathbb{N}\setminus S)_{*})$ we add the elements of $W_{k}$ into the digit sequence $(a_{n}(y))_{n=1}^{\infty}$ at position $M_{k}$ to define a new sequence. Then $F_{S}$ is the set of points whose regular continued fraction expansions are given by these new sequences. It is clear from the construction that for any $x\in F_{S}$, the digit sequence $\{a_{n}(x)\}_{n\in\mathbb{N}}$ is not monotone increasing.
@@
$\dim_{\rm H}F_{S}\geq\dim_{\rm H}R_{t,L}((\mathbb{N}\setminus S)_{*})=\frac{1}{2}.$
@@
$\bigcup_{n\in\mathbb{N}}\bigcap_{x\in F_{S}}\{a_{n}(x)\}\cap W_{k}=W_{k}.$
@@
$\begin{split}d_{\rm B}\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in F_{S}}\{a_{n}(x)\}\cap S\right)&\geq\limsup_{k\to\infty}\frac{\#\left(\bigcup_{n\in\mathbb{N}}\bigcap_{x\in F_{S}}\{a_{n}(x)\}\cap W_{k}\right)}{\#(\mathbb{N}\cap I_{k})}\\ &=\lim_{k\to\infty}\frac{\#W_{k}}{\#(\mathbb{N}\cap I_{k})}=d_{\rm B}(S).\end{split}$